Tuesday, July 11, 2006
- Temporary repairs to the Dental School are continuing.
- The hot water heaters are in place and being connected to the building plumbing.
- The desiccant air dryer for medical air has been installed.
- Some of the old black iron pipe is being replaced with copper to meet current codes.
- Sump pumps are in place and connected.
- The environmental remediation at 1542 Tulane also continues to progress.
- The demolition of the flood damaged walls in the basement continues. We estimate that the work is about 90% complete.
- Insulation in one of the basement walls being demolished is being tested for asbestos.
- Following the installation of new floor tile in the Residence Hall, the contractor is completing the installation of new door stops and thresholds.
- The installation of new carpeting in the two large Lecture Halls at the MEB has begun.
- Floor preparations for the application of the new epoxy surface continued today in the MEB Gross Lab and Morgue.
- The installation of replacement ceiling tile in the MDL is approximately 50% complete.
- The contractor has begun the installation of new projector supports in the two large Lecture Halls at the MEB.